find vertex, axis of symetry and max/min value of the parabola by transforming into vertex form
f(x)=3x^2-12x+2
and the selection bar for choosing the topic wasnt working correctley thats why i didnt choose one sorry

Answer by lwsshak3(11628)

find vertex, axis of symetry and max/min value of the parabola by transforming into vertex form
f(x)=3x^2-12x+2
complete the square:
f(x)=3(x^2-4x+4)-12+2
f(x)=3(x-2)^2-10
This is an equation of a parabola that opens upward:
Its standard form: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex, A is a coefficient that affects the slope or steepness of the curve.
For given parabola:
vertex: (2,-10)
axis of symmetry: x=2
minimum value: -10